# A Unified Linear Programming Framework for Offline Reward Learning from Human Demonstrations and Feedback

### v1

@@ 5 Numerical Experiments
The left side of Figure 1 compares the errors associated with each IRL algorithm. The results indicate that our LP-based algorithms generally outperform the bi-level optimization-based MLE algorithm [18], demonstrating that LP is more sample-efficient in addressing ambiguity in the dynamics and the expert policy. The solution set with a smaller relaxation level $\epsilon_{g}=0.001/\sqrt{N}$ (LP-IRL-2) exhibits better performance than that with a greater relaxation level $\epsilon_{g}=0.01/\sqrt{N}$ (LP-IRL-1). This is consistent with the optimality-feasibility trade-off discussed in Section 3.2. Additionally, the integration of IRL and RLHF data leads to improved performance, as predicted. The use of continuous feedback (LP-IRL-C) is even more effective than discrete feedback (LP-IRL-D) by facilitating stricter constraints.
@@
In numerical experiments for offline RLHF, the human feedback data is generated following the greedy model. The right side of Figure 1 compares the reward function obtained from LP (4.3) and the pessimistic MLE algorithm proposed by [19] under the BTL model. In the LP algorithm, the error decreases rapidly as the number of samples increases, whereas the error in the MLE algorithm decreases more slowly. This result is consistent with the discussion in Appendix F, suggesting that the MLE algorithm might be inefficient or even fail if the human evaluator deviates from the assumed model, whereas LP does not.
@@ Appendix A Related Work
One of the foundational works in IRL [12] introduced the concept of characterizing a set of reward functions for which a given policy is optimal using an LP formulation. This idea has been further developed in subsequent literature [22, 24, 23, 25], as outlined in the introduction. Recently proposed practical offline imitation learning (IL) algorithms, including ValueDICE [40], IQ-Learn [41], OPIRL [42], and ReCOIL [43], address an occupancy matching problem that minimizes the statistical divergence between the learner and the expert distribution. These algorithms exploit the duality of the LP formulation to obtain tractable algorithms, as extensively discussed in [43]. Despite the practical success of these algorithms, the resulting reward function and policy depend on the model in use, and they lack theoretical performance guarantees, such as provable sample efficiency.
@@
In offline RLHF, we impose a margin-based constraint on the solution set, which allows for the alignment of reward functions with preference data without assuming any preference models of human evaluators. The concept of employing a margin constraint originated in the early imitation learning literature. Specifically, maximum margin planning (MMP) [44, 45] estimates the reward function such that the expert policy achieves a higher expected reward than all other policies by imposing a margin constraint in the reward optimization problem. Recently, [46] introduced Rank-Game, a two-player game formulation between a policy agent, which optimizes the policy given a reward function, and a reward agent, which aligns the reward function with offline pairwise preference data. Their algorithm is model-free, as the reward agent minimizes ranking loss without relying on a specific preference model. A unification of demonstration and preference data is also proposed in their work, similar to our approach in the LP framework.
Appendix B Proof of Lemma 2
@@
$\mathbb{P}(X^{\top}(K_{\mathcal{D}}w_{\tilde{d}}-(1-\gamma)\mu_{0})\leq\epsilon_{x})\geq 1-\delta^{\prime},$ (B.1)
@@
$x_{i}^{\top}(K_{H}-K)w_{\tilde{d}}\leq(1+\gamma)\gamma^{H}B$ (B.2)
@@
$x_{i}^{\top}(K_{\mathcal{D}}-K_{H})w_{\tilde{d}}\geq\epsilon_{xi}-(1+\gamma)\gamma^{H}B$ (B.3)
@@
$d_{e}^{H}(s,a):=(1-\gamma)\sum_{h=0}^{H-1}\gamma^{h}P^{\pi_{e}}_{\mu_{0}}(s_{h}=s,a_{h}=a)\quad\forall(s,a)\in S\times A.$ (B.4)
@@
$d_{e}^{P,H}(s,a,s^{\prime}):=(1-\gamma)\sum_{h=0}^{H-1}\gamma^{h}P^{\pi_{e}}_{\mu_{0}}(s_{h}=s,a_{h}=a,s_{h+1}=s^{\prime})\quad\forall(s,a,s^{\prime})\in S\times A\times S.$ (B.5)
@@
$K_{H}(s^{\prime},(s,a)):=d_{e}^{H}(s,a)\mathbf{1}_{\{s=s^{\prime}\}}-\gamma d_{e}^{P,H}(s,a,s^{\prime})\quad\forall(s,a,s^{\prime})\in S\times A\times S.$ (B.6)
@@
$\begin{split}|x_{i}^{\top}(K-K_{H})w_{\tilde{d}}|&\leq\sum_{s^{\prime}\in S}|x_{i}(s^{\prime})|\sum_{h=H}^{\infty}(1-\gamma)\gamma^{h}\sum_{a\in A}\mathbb{P}^{\pi_{e}}_{\mu_{0}}(s_{h}=s^{\prime},a_{h}=a)w_{\tilde{d}}(s^{\prime},a)\\ &+\sum_{s^{\prime}\in S}|x_{i}(s^{\prime})|\sum_{h=H}^{\infty}(1-\gamma)\gamma^{h+1}\sum_{(s,a)\in S\times A}\mathbb{P}^{\pi_{e}}_{\mu_{0}}(s_{h}=s,a_{h}=a,s_{h+1}=s^{\prime})w_{\tilde{d}}(s,a)\\ &\leq\sum_{s^{\prime}\in S}\sum_{h=H}^{\infty}(1-\gamma)\gamma^{h}\mathbb{P}^{\pi_{e}}_{\mu_{0}}(s_{h}=s^{\prime})B+\sum_{s^{\prime}\in S}\sum_{h=H}^{\infty}(1-\gamma)\gamma^{h+1}\mathbb{P}^{\pi_{e}}_{\mu_{0}}(s_{h+1}=s^{\prime})B\\ &=\gamma^{H}B+\gamma^{H+1}B=(1+\gamma)\gamma^{H}B.\end{split}$ (B.7)
@@
$\begin{split}z(\tau):=&\sum_{s^{\prime}\in S}x_{i}(s^{\prime})\sum_{(s,a)\in S\times A}w_{\tilde{d}}(s,a)(1-\gamma)\times\sum_{h=0}^{H-1}\left[\gamma^{h}\mathbf{1}_{\{s_{h}=s,a_{h}=a\}}-\gamma^{h+1}\mathbf{1}_{\{s_{h}=s,a_{h}=a,s_{h+1}=s^{\prime}\}}\right].\end{split}$ (B.8)
@@
$\begin{split}|z(\tau)|&\leq B(1-\gamma)\sum_{s^{\prime}\in S}\sum_{(s,a)\in S\times A}\sum_{h=0}^{H-1}\left[\gamma^{h}\mathbf{1}_{\{s_{h}=s,a_{h}=a\}}+\gamma^{h+1}\mathbf{1}_{\{s_{h}=s,a_{h}=a,s_{h+1}=s^{\prime}\}}\right]\\ &\leq B(1-\gamma)\left(\frac{1-\gamma^{H}}{1-\gamma}+\frac{\gamma-\gamma^{H+1}}{1-\gamma}\right)=B(1+\gamma)(1-\gamma^{H}),\end{split}$ (B.9)
@@
$\mathbb{P}\left(x_{i}^{\top}(K_{\mathcal{D}}-K_{H})w_{\tilde{d}}\geq\epsilon\right)\leq\exp\left(-\frac{N\epsilon^{2}}{2B^{2}(1+\gamma)^{2}(1-\gamma^{H})^{2}}\right)\quad\forall\epsilon\geq 0.$ (B.10)
@@
$\mathbb{P}\left(x_{i}^{\top}(K_{\mathcal{D}}-K_{H})w_{\tilde{d}}\geq\epsilon_{xi}-(1+\gamma)\gamma^{H}B\right)\leq\frac{\delta^{\prime}}{N_{x}}.$ (B.11)
@@
$\mathbb{P}\left(x_{i}^{\top}(K_{\mathcal{D}}-K_{H})w_{\tilde{d}}+x_{i}^{\top}(K_{H}-K)w_{\tilde{d}}\geq\epsilon_{xi}\right)\leq\mathbb{P}\left(x_{i}^{\top}(K_{\mathcal{D}}-K_{H})w_{\tilde{d}}+(1+\gamma)\gamma^{H}B\geq\epsilon_{xi}\right)\leq\frac{\delta^{\prime}}{N_{x}}.$ (B.12)
@@
$\mathbb{P}\left(X^{\top}(K_{\mathcal{D}}-K)w_{\tilde{d}}\geq\epsilon_{x}\right)\leq\delta^{\prime}.$ (B.13)
@@
$\mathbb{P}\left(X^{\top}(K_{\mathcal{D}}w_{\tilde{d}}-(1-\gamma)\mu_{0})\geq\epsilon_{x}\right)\leq\delta^{\prime},$ (B.14)
@@
Appendix C Proof of Theorem 1
The proof is comprised of three steps. In the first step, we employ a concentration bound to establish a limit on the difference between $u^{\top}\mathbf{1}$ and $u_{\mathcal{D}}^{\top}\mathbf{1}$, ensuring that $\mathbb{P}(u^{\top}\mathbf{1}-u_{\mathcal{D}}^{\top}\mathbf{1}\geq-\epsilon_{u1})\geq 1-\delta$ for a certain $\epsilon_{u1}$. Next, from the feasibility of $w_{\tilde{d}}$ proven by Lemma 2 and using the optimality conditions in (3.11), we can show that $\mathbb{P}(u_{\mathcal{D}}^{\top}\mathbf{1}\geq u_{\mathcal{D}}^{\top}w_{\tilde{d}}-\epsilon_{g},\;\forall\tilde{d}\in D_{B})\geq 1-\delta$ holds. The final step is to bound the difference between $u_{\mathcal{D}}^{\top}w_{\tilde{d}}$ and $u^{\top}w_{\tilde{d}}$, showing that $\mathbb{P}(u_{\mathcal{D}}^{\top}w_{\tilde{d}}-u^{\top}w_{\tilde{d}}\geq-\epsilon_{u2},\;\forall\tilde{d}\in D_{B})\geq 1-\delta$ for a specific $\epsilon_{u2}$. Combining these three results with the union bound completes the proof.
@@ Appendix C Proof of Theorem 1
$\begin{split}|r^{\top}(d_{e}-d_{e}^{H})|&\leq(1-\gamma)\sum_{(s,a)\in S\times A}|r(s,a)|\sum_{h=H}^{\infty}\gamma^{h}P^{\pi}_{\mu_{0}}(s_{h}=s,a_{h}=a)\\ &\leq(1-\gamma)\sum_{(s,a)\in S\times A}\sum_{h=H}^{\infty}\gamma^{h}P^{\pi}_{\mu_{0}}(s_{h}=s,a_{h}=a)=\gamma^{H}.\end{split}$ (C.1)
@@
$z^{\prime}(\tau):=(1-\gamma)\sum_{(s,a)\in S\times A}r(s,a)\sum_{h=0}^{H-1}\gamma^{h}\mathbf{1}_{\{s_{h}=s,a_{h}=a\}},$ (C.2)
@@
$\mathbb{P}(r^{\top}(d_{e}^{H}-\hat{d}_{e})\leq-\epsilon)\leq\exp\left(\frac{-N\epsilon^{2}}{2(1-\gamma^{H})^{2}}\right)\quad\forall\epsilon\geq 0.$ (C.3)
@@
$\begin{split}\mathbb{P}(u^{\top}\mathbf{1}-u_{\mathcal{D}}^{\top}\mathbf{1}\leq-\epsilon_{u1})&=\mathbb{P}(r^{\top}(d_{e}^{H}-\hat{d}_{e})+r^{\top}(d_{e}-d_{H})\leq-\epsilon_{u1})\\ &\leq\mathbb{P}(r^{\top}(d_{e}^{H}-\hat{d}_{e})\leq-\epsilon_{u1}+\gamma^{H})\leq\delta.\end{split}$ (C.4)
Next, by Lemma 2, $w_{\tilde{d}}$ is a feasible solution to (3.8) with probability at least $1-\frac{\delta}{2^{|S||A|}}$. If $w_{\tilde{d}}$ is a feasible solution to (3.8), then the following inequalities hold:
$\begin{split}u_{\mathcal{D}}^{\top}\mathbf{1}&\overset{(i)}{\geq}(1-\gamma)\mu_{0}^{\top}Xv_{\mathcal{D}}+\epsilon_{x}^{\top}v_{\mathcal{D}}-\epsilon_{g}\\ &\overset{(ii)}{\geq}w_{\tilde{d}}^{\top}K_{\mathcal{D}}^{\top}Xv_{\mathcal{D}}-\epsilon_{g}\\ &\overset{(iii)}{\geq}w_{\tilde{d}}^{\top}u_{\mathcal{D}}-\epsilon_{g}\\ \end{split}$ (C.5)
The inequality $(i)$ holds by the duality gap constraint, $(ii)$ holds because $w_{\tilde{d}}$ is feasible to (3.8), and $(iii)$ holds by the feasibility constraint $K_{\mathcal{D}}^{\top}Xv_{\mathcal{D}}\geq u_{\mathcal{D}}$ and $w_{\tilde{d}}\geq 0$. Therefore, we get
$\mathbb{P}(u_{\mathcal{D}}^{\top}\mathbf{1}\geq u_{\mathcal{D}}^{\top}w_{\tilde{d}}-\epsilon_{g})\geq 1-\frac{\delta}{2^{|S||A|}}.$ (C.6)
@@
$\mathbb{P}(u_{\mathcal{D}}^{\top}\mathbf{1}\geq u_{\mathcal{D}}^{\top}w_{\tilde{d}}-\epsilon_{g},\;\forall\tilde{d}\in D_{B})\geq 1-\delta.$ (C.7)
@@
$\mathbb{P}(u_{\mathcal{D}}^{\top}w_{\tilde{d}}-u^{\top}w_{\tilde{d}}\geq-\epsilon_{u2})\geq 1-\frac{\delta}{2^{|S||A|}},$ (C.8)
@@
$\mathbb{P}(u_{\mathcal{D}}^{\top}w_{\tilde{d}}-u^{\top}w_{\tilde{d}}\geq-\epsilon_{u2},\;\forall\tilde{d}\in D_{B})\geq 1-\delta,$ (C.9)
Taking union bound to the above three cases and using $u^{\top}\mathbf{1}=r^{\top}d_{e}$ and $u^{\top}w_{\tilde{d}}=r^{\top}\tilde{d}$ from Lemma 1, the conclusion holds as
$\mathbb{P}(r^{\top}d_{e}\geq r^{\top}\tilde{d}-\epsilon,\;\forall\tilde{d}\in D_{B})=\mathbb{P}(u^{\top}\mathbf{1}\geq u^{\top}w_{\tilde{d}}-\epsilon_{g}-\epsilon_{u1}-\epsilon_{u2},\;\forall\tilde{d}\in D_{B})\geq 1-3\delta.$ (C.10)

### v3

@@ 5 Numerical Experiments
The left side of Figure 1 compares the errors associated with each IRL algorithm. The results indicate that our LP-based algorithms generally outperform the bi-level optimization-based MLE algorithm (Zeng et al., 2023), demonstrating that LP is more sample-efficient in addressing ambiguity in the dynamics and the expert policy. The solution set with a smaller relaxation level $\epsilon_{g}=0.001/\sqrt{N}$ (LP-IRL-2) exhibits better performance than that with a greater relaxation level $\epsilon_{g}=0.01/\sqrt{N}$ (LP-IRL-1). This is consistent with the optimality-feasibility trade-off discussed in Section 3.2. Additionally, the integration of IRL and RLHF data leads to improved performance, as predicted. The use of continuous feedback (LP-IRL-C) is even more effective than discrete feedback (LP-IRL-D) by facilitating stricter constraints.
@@
In numerical experiments for offline RLHF, the human feedback data is generated following the greedy model. The right side of Figure 1 compares the reward function obtained from LP (29) and the pessimistic MLE algorithm proposed by (Zhu et al., 2023) under the BTL model. In the LP algorithm, the error decreases rapidly as the number of samples increases, whereas the error in the MLE algorithm decreases more slowly. This result is consistent with the discussion in Appendix F, suggesting that the MLE algorithm might be inefficient or even fail if the human evaluator deviates from the assumed model, whereas LP does not.
@@ 6 Conclusion
Acknowledgments
The authors would like to express their gratitude for the support that made this work possible. J. Zhang was supported by the MIT School of Engineering Postdoctoral Fellowship Program for Engineering Excellence. A. Ozdaglar acknowledges support from the AI Accelerator program.
Impact Statement
This paper aims to advance the understanding and efficiency of reward learning. We believe our work has the potential to benefit sequential decision-making systems in fields such as robotics, autonomous driving, and language models, though no immediate specific societal concerns are directly raised by our findings.
@@ Appendix A Related Work
One of the foundational works in IRL (Ng & Russell, 2000) introduced the concept of characterizing a set of reward functions for which a given policy is optimal using an LP formulation. This idea has been further developed in subsequent literature (Metelli et al., 2021; Lindner et al., 2022; Metelli et al., 2023; Zhao et al., 2023), as outlined in the introduction. Recently proposed practical offline imitation learning (IL) algorithms, including ValueDICE (Kostrikov et al., 2019), IQ-Learn (Garg et al., 2021), OPIRL (Hoshino et al., 2022), and ReCOIL (Sikchi et al., 2023b), address an occupancy matching problem that minimizes the statistical divergence between the learner and the expert distribution. These algorithms exploit the duality of the LP formulation to obtain tractable algorithms, as extensively discussed in (Sikchi et al., 2023b). Despite the practical success of these algorithms, the resulting reward function and policy depend on the model in use, and they lack theoretical performance guarantees, such as provable sample efficiency.
@@
In offline RLHF, we impose a margin-based constraint on the solution set, which allows for the alignment of reward functions with preference data without assuming any preference models of human evaluators. The concept of employing a margin constraint originated in the early imitation learning literature. Specifically, maximum margin planning (MMP) (Ratliff et al., 2006, 2009) estimates the reward function such that the expert policy achieves a higher expected reward than all other policies by imposing a margin constraint in the reward optimization problem. Recently, (Sikchi et al., 2023a) introduced Rank-Game, a two-player game formulation between a policy agent, which optimizes the policy given a reward function, and a reward agent, which aligns the reward function with offline pairwise preference data. Their algorithm is model-free, as the reward agent minimizes ranking loss without relying on a specific preference model. A unification of demonstration and preference data is also proposed in their work, similar to our approach in the LP framework.
Appendix B Proof of Lemma 3.2
@@
$\mathbb{P}(X^{\top}(K_{\mathcal{D}}w_{\tilde{d}}-(1-\gamma)\mu_{0})\leq\epsilon_{x})\geq 1-\delta^{\prime},$ (32)
@@
$x_{i}^{\top}(K_{H}-K)w_{\tilde{d}}\leq(1+\gamma)\gamma^{H}B$ (33)
@@
$x_{i}^{\top}(K_{\mathcal{D}}-K_{H})w_{\tilde{d}}\geq\epsilon_{xi}-(1+\gamma)\gamma^{H}B$ (34)
@@
$d_{e}^{H}(s,a):=(1-\gamma)\sum_{h=0}^{H-1}\gamma^{h}P^{\pi_{e}}_{\mu_{0}}(s_{h}=s,a_{h}=a)\quad\forall(s,a)\in S\times A.$ (35)
@@
$d_{e}^{P,H}(s,a,s^{\prime}):=(1-\gamma)\sum_{h=0}^{H-1}\gamma^{h}P^{\pi_{e}}_{\mu_{0}}(s_{h}=s,a_{h}=a,s_{h+1}=s^{\prime})\quad\forall(s,a,s^{\prime})\in S\times A\times S.$ (36)
@@
$K_{H}(s^{\prime},(s,a)):=d_{e}^{H}(s,a)\mathbf{1}_{\{s=s^{\prime}\}}-\gamma d_{e}^{P,H}(s,a,s^{\prime})\quad\forall(s,a,s^{\prime})\in S\times A\times S.$ (37)
@@
$\begin{split}|x_{i}^{\top}(K-K_{H})w_{\tilde{d}}|&\leq\sum_{s^{\prime}\in S}|x_{i}(s^{\prime})|\sum_{h=H}^{\infty}(1-\gamma)\gamma^{h}\sum_{a\in A}\mathbb{P}^{\pi_{e}}_{\mu_{0}}(s_{h}=s^{\prime},a_{h}=a)w_{\tilde{d}}(s^{\prime},a)\\ &+\sum_{s^{\prime}\in S}|x_{i}(s^{\prime})|\sum_{h=H}^{\infty}(1-\gamma)\gamma^{h+1}\sum_{(s,a)\in S\times A}\mathbb{P}^{\pi_{e}}_{\mu_{0}}(s_{h}=s,a_{h}=a,s_{h+1}=s^{\prime})w_{\tilde{d}}(s,a)\\ &\leq\sum_{s^{\prime}\in S}\sum_{h=H}^{\infty}(1-\gamma)\gamma^{h}\mathbb{P}^{\pi_{e}}_{\mu_{0}}(s_{h}=s^{\prime})B+\sum_{s^{\prime}\in S}\sum_{h=H}^{\infty}(1-\gamma)\gamma^{h+1}\mathbb{P}^{\pi_{e}}_{\mu_{0}}(s_{h+1}=s^{\prime})B\\ &=\gamma^{H}B+\gamma^{H+1}B=(1+\gamma)\gamma^{H}B.\end{split}$ (38)
@@
$\begin{split}z(\tau):=&\sum_{s^{\prime}\in S}x_{i}(s^{\prime})\sum_{(s,a)\in S\times A}w_{\tilde{d}}(s,a)(1-\gamma)\times\sum_{h=0}^{H-1}\left[\gamma^{h}\mathbf{1}_{\{s_{h}=s,a_{h}=a\}}-\gamma^{h+1}\mathbf{1}_{\{s_{h}=s,a_{h}=a,s_{h+1}=s^{\prime}\}}\right].\end{split}$ (39)
@@
$\begin{split}|z(\tau)|&\leq B(1-\gamma)\sum_{s^{\prime}\in S}\sum_{(s,a)\in S\times A}\sum_{h=0}^{H-1}\left[\gamma^{h}\mathbf{1}_{\{s_{h}=s,a_{h}=a\}}+\gamma^{h+1}\mathbf{1}_{\{s_{h}=s,a_{h}=a,s_{h+1}=s^{\prime}\}}\right]\\ &\leq B(1-\gamma)\left(\frac{1-\gamma^{H}}{1-\gamma}+\frac{\gamma-\gamma^{H+1}}{1-\gamma}\right)=B(1+\gamma)(1-\gamma^{H}),\end{split}$ (40)
@@
$\mathbb{P}\left(x_{i}^{\top}(K_{\mathcal{D}}-K_{H})w_{\tilde{d}}\geq\epsilon\right)\leq\exp\left(-\frac{N\epsilon^{2}}{2B^{2}(1+\gamma)^{2}(1-\gamma^{H})^{2}}\right)\quad\forall\epsilon\geq 0.$ (41)
@@
$\mathbb{P}\left(x_{i}^{\top}(K_{\mathcal{D}}-K_{H})w_{\tilde{d}}\geq\epsilon_{xi}-(1+\gamma)\gamma^{H}B\right)\leq\frac{\delta^{\prime}}{N_{x}}.$ (42)
@@
$\mathbb{P}\left(x_{i}^{\top}(K_{\mathcal{D}}-K_{H})w_{\tilde{d}}+x_{i}^{\top}(K_{H}-K)w_{\tilde{d}}\geq\epsilon_{xi}\right)\leq\mathbb{P}\left(x_{i}^{\top}(K_{\mathcal{D}}-K_{H})w_{\tilde{d}}+(1+\gamma)\gamma^{H}B\geq\epsilon_{xi}\right)\leq\frac{\delta^{\prime}}{N_{x}}.$ (43)
@@
$\mathbb{P}\left(X^{\top}(K_{\mathcal{D}}-K)w_{\tilde{d}}\geq\epsilon_{x}\right)\leq\delta^{\prime}.$ (44)
@@
$\mathbb{P}\left(X^{\top}(K_{\mathcal{D}}w_{\tilde{d}}-(1-\gamma)\mu_{0})\geq\epsilon_{x}\right)\leq\delta^{\prime},$ (45)
@@
Appendix C Proof of Theorem 3.3
The proof is comprised of three steps. In the first step, we employ a concentration bound to establish a limit on the difference between $u^{\top}\mathbf{1}$ and $u_{\mathcal{D}}^{\top}\mathbf{1}$, ensuring that $\mathbb{P}(u^{\top}\mathbf{1}-u_{\mathcal{D}}^{\top}\mathbf{1}\geq-\epsilon_{u1})\geq 1-\delta$ for a certain $\epsilon_{u1}$. Next, from the feasibility of $w_{\tilde{d}}$ proven by Lemma 3.2 and using the optimality conditions in (17), we can show that $\mathbb{P}(u_{\mathcal{D}}^{\top}\mathbf{1}\geq u_{\mathcal{D}}^{\top}w_{\tilde{d}}-\epsilon_{g},\;\forall\tilde{d}\in D_{B})\geq 1-\delta$ holds. The final step is to bound the difference between $u_{\mathcal{D}}^{\top}w_{\tilde{d}}$ and $u^{\top}w_{\tilde{d}}$, showing that $\mathbb{P}(u_{\mathcal{D}}^{\top}w_{\tilde{d}}-u^{\top}w_{\tilde{d}}\geq-\epsilon_{u2},\;\forall\tilde{d}\in D_{B})\geq 1-\delta$ for a specific $\epsilon_{u2}$. Combining these three results with the union bound completes the proof.
@@
$\begin{split}|r^{\top}(d_{e}-d_{e}^{H})|&\leq(1-\gamma)\sum_{(s,a)\in S\times A}|r(s,a)|\sum_{h=H}^{\infty}\gamma^{h}P^{\pi}_{\mu_{0}}(s_{h}=s,a_{h}=a)\\ &\leq(1-\gamma)\sum_{(s,a)\in S\times A}\sum_{h=H}^{\infty}\gamma^{h}P^{\pi}_{\mu_{0}}(s_{h}=s,a_{h}=a)=\gamma^{H}.\end{split}$ (46)
@@
$z^{\prime}(\tau):=(1-\gamma)\sum_{(s,a)\in S\times A}r(s,a)\sum_{h=0}^{H-1}\gamma^{h}\mathbf{1}_{\{s_{h}=s,a_{h}=a\}},$ (47)
@@
$\mathbb{P}(r^{\top}(d_{e}^{H}-\hat{d}_{e})\leq-\epsilon)\leq\exp\left(\frac{-N\epsilon^{2}}{2(1-\gamma^{H})^{2}}\right)\quad\forall\epsilon\geq 0.$ (48)
@@
$\begin{split}\mathbb{P}(u^{\top}\mathbf{1}-u_{\mathcal{D}}^{\top}\mathbf{1}\leq-\epsilon_{u1})&=\mathbb{P}(r^{\top}(d_{e}^{H}-\hat{d}_{e})+r^{\top}(d_{e}-d_{H})\leq-\epsilon_{u1})\\ &\leq\mathbb{P}(r^{\top}(d_{e}^{H}-\hat{d}_{e})\leq-\epsilon_{u1}+\gamma^{H})\leq\delta.\end{split}$ (49)
Next, by Lemma 3.2, $w_{\tilde{d}}$ is a feasible solution to (14) with probability at least $1-\frac{\delta}{2^{|S||A|}}$. If $w_{\tilde{d}}$ is a feasible solution to (14), then the following inequalities hold:
$\begin{split}u_{\mathcal{D}}^{\top}\mathbf{1}&\overset{(i)}{\geq}(1-\gamma)\mu_{0}^{\top}Xv_{\mathcal{D}}+\epsilon_{x}^{\top}v_{\mathcal{D}}-\epsilon_{g}\\ &\overset{(ii)}{\geq}w_{\tilde{d}}^{\top}K_{\mathcal{D}}^{\top}Xv_{\mathcal{D}}-\epsilon_{g}\\ &\overset{(iii)}{\geq}w_{\tilde{d}}^{\top}u_{\mathcal{D}}-\epsilon_{g}\\ \end{split}$ (50)
The inequality $(i)$ holds by the duality gap constraint, $(ii)$ holds because $w_{\tilde{d}}$ is feasible to (14), and $(iii)$ holds by the feasibility constraint $K_{\mathcal{D}}^{\top}Xv_{\mathcal{D}}\geq u_{\mathcal{D}}$ and $w_{\tilde{d}}\geq 0$. Therefore, we get
$\mathbb{P}(u_{\mathcal{D}}^{\top}\mathbf{1}\geq u_{\mathcal{D}}^{\top}w_{\tilde{d}}-\epsilon_{g})\geq 1-\frac{\delta}{2^{|S||A|}}.$ (51)
@@
$\mathbb{P}(u_{\mathcal{D}}^{\top}\mathbf{1}\geq u_{\mathcal{D}}^{\top}w_{\tilde{d}}-\epsilon_{g},\;\forall\tilde{d}\in D_{B})\geq 1-\delta.$ (52)
@@
$\mathbb{P}(u_{\mathcal{D}}^{\top}w_{\tilde{d}}-u^{\top}w_{\tilde{d}}\geq-\epsilon_{u2})\geq 1-\frac{\delta}{2^{|S||A|}},$ (53)
@@
$\mathbb{P}(u_{\mathcal{D}}^{\top}w_{\tilde{d}}-u^{\top}w_{\tilde{d}}\geq-\epsilon_{u2},\;\forall\tilde{d}\in D_{B})\geq 1-\delta,$ (54)
Taking union bound to the above three cases and using $u^{\top}\mathbf{1}=r^{\top}d_{e}$ and $u^{\top}w_{\tilde{d}}=r^{\top}\tilde{d}$ from Lemma 3.1, the conclusion holds as
$\mathbb{P}(r^{\top}d_{e}\geq r^{\top}\tilde{d}-\epsilon,\;\forall\tilde{d}\in D_{B})=\mathbb{P}(u^{\top}\mathbf{1}\geq u^{\top}w_{\tilde{d}}-\epsilon_{g}-\epsilon_{u1}-\epsilon_{u2},\;\forall\tilde{d}\in D_{B})\geq 1-3\delta.$ (55)